OpenStudy (vianne):

Holly missed 6 of 185 school days this year. Approximately what percent of school days was Holly in attendance?

OpenStudy (vianne):

@waterineyes

OpenStudy (vianne):

69% 3% 97% 31% Those are the answer choices..

OpenStudy (vianne):

Don't give me just the answer, but explain how you solved the problem so that I can understand it.

OpenStudy (anonymous):

That is incorrect

OpenStudy (anonymous):

See, this is very easy..

OpenStudy (anonymous):

If Holly misses 6 days, then how many days he or she attended?

OpenStudy (vianne):

179

OpenStudy (anonymous):

185 - 6 = 179 179/185 = ~97%

OpenStudy (vianne):

Thanks! It's correct, and thanks for explaining how you solved it. :)

OpenStudy (anonymous):

Good..

OpenStudy (vianne):

Martha estimated there were 84 marbles in a jar for a contest. The actual number of marbles in the jar was 106. What was the percent error of Martha's estimation? Would it be 22%?

OpenStudy (anonymous):

So, percentage is given by : \[\% \ attendence = \frac{Attended \ days}{Total \ Days} \times 100\]

OpenStudy (vianne):

Umm, would it?

OpenStudy (anonymous):

What?

OpenStudy (vianne):

.................. I got it wrong.. It was actually 20.75% D:

OpenStudy (anonymous):

Which one??

OpenStudy (anonymous):

Since you have posted two questions?? Which one?? First one or second one?

OpenStudy (vianne):

Erm.. I need help with this one...... Nathan took his car in for service and repairs. He had a coupon for 10% off. Original Prices • Labor $242.30 • Parts $324.53 • Other $66.02 Which amount is closest to Nathan's total costs after the 10% discount and including 7% sales tax? [Assume tax is applied after the discount is applied.]

OpenStudy (anonymous):

Second one most likely

OpenStudy (anonymous):

Sorry, you guys carry on.. OS is not upto the mark for me today. :(

OpenStudy (vianne):

:(

OpenStudy (anonymous):

The second one would be: 106-84 = 22 22/106 = 20.75%

OpenStudy (vianne):

Oh, what about the third one?

OpenStudy (anonymous):

Labor -$242.30 $242.30 x .10 = $24.23 discount $242.30 - $24.23 = 218.07 after discount and before tax $218.07 x 0.07 = $15.26 + $218.07 = $233.33 after discount and tax Parts - $324.53 $324.53 x .10 = $35.45 discount $324.53 - $35.45 = $289.08 after discount and before tax $289.08 x 0.07 = $20.24 + $289.08 = $309.32 after discount and tax Other - $66.02 $66.02 x .10 = $6.60 discount $66.02 - $6.60 =$59.42 after discount and before tax $59.42 x 0.07 = $4.16 + $59.42 = $63.58 after discount and tax

OpenStudy (anonymous):

I don't know if that is how you wanted the answer..

OpenStudy (vianne):

I found the answer.

OpenStudy (anonymous):

Did it want you to add them all first. Apply the discount then tax?